OpenStudy (anonymous):

Help, what am I doing wrong? Jessie and Bob are financing $425,500 to purchase a house. They obtained a 30/8 balloon mortgage at 6.55%. What will their balloon payment be?

OpenStudy (anonymous):

i think u should multiple rate in whole the mortgage

OpenStudy (anonymous):

i mean : 178800*(3.95/100)

OpenStudy (anonymous):

I solved and got 379675.59529504773077296

OpenStudy (anonymous):

It is close to my answer choices of $379,659.89 $637,130.99 $357,491.39 $380,040.19

OpenStudy (anonymous):

but I don't know which to chose.

OpenStudy (anonymous):

Is it 380,040.19 if i rounded my answer?

OpenStudy (anonymous):

formula is Bm=P{(1+i)^nt - (1+i)^m / (1+i)^nt - 1} I=0.0655/12 n=12 t=30 nt=360 P=425500

OpenStudy (anonymous):

I got 377598.80

OpenStudy (anonymous):

sri , but i don't know this formula !!!

OpenStudy (anonymous):

now I need to add the 0.0655/12 + my 377598.80 =

OpenStudy (anonymous):

207679.34 + 377598.80 = and this time now I got 585278.14

OpenStudy (anonymous):

What am I doing wrong?

OpenStudy (anonymous):

woops I wrote the wrong question

OpenStudy (anonymous):

Jessie and Bob are financing $425,500 to purchase a house. They obtained a 30/8 balloon mortgage at 6.55%. What will their balloon payment be?

OpenStudy (anonymous):

ok dear let find out the question step by step . u said balloon mortgage . it's mean if they didn't pay their loan the must pay more ? and the rate of money the have to pay is 15/7 am i right ?

OpenStudy (anonymous):

The 15/7 mean 15= Initial interest period 7 means interest rate adjustment

OpenStudy (anonymous):

I have a good grade in this class, but this mortgage and balloon stuff is killing me

OpenStudy (anonymous):

i think it's so hard . i try to find out the solution . u know english is not my mother tongue so i should check for some meaning to find out the question .

OpenStudy (anonymous):

m=96

OpenStudy (anonymous):

would u like to describe the formula's parameter for me ?

OpenStudy (anonymous):

What is that? I am going to take a picture of the formula with the numbers plugged in them for you

OpenStudy (anonymous):

i read it but i want to know how you find out the numbers ?

OpenStudy (anonymous):

I just plugged it into my scientific calculator

OpenStudy (anonymous):

i'm so sri but i think i can't help you in this problem .

OpenStudy (anonymous):

i should say apologize for taking ur time . i'm really sri

OpenStudy (anonymous):